# EMMA: End-to-End Multimodal Model for Autonomous Driving

### v2

@@ 3.1 End-to-End Motion Planning on Public Datasets
We conduct the end-to-end planner trajectory generation experiments on two public datasets, the Waymo Open Motion Dataset (WOMD) (Chen et al., 2024a) and the nuScenes dataset (Caesar et al., 2020). EMMA is trained with the simplest end-to-end planner trajectory generation formulation as in Equation 2. That is, given camera images, ego vehicle history, and driving intent, the model is asked to predict the future ego waypoints for a certain time horizon.
Method ADE 1s ADE 3s ADE 5s ADE 8s
MotionLM (Seff et al., 2023) 0.045 0.251 0.694 1.766
Wayformer* (Nayakanti et al., 2023) 0.044 0.250 0.640 1.517
EMMA 0.032 0.248 0.681 1.718
EMMA+ 0.030 0.225 0.610 1.553
Table 1: End-to-end motion planning experiments on WOMD planning benchmark (Chen et al., 2024a). EMMA+ denotes the models are pre-trained on our internal data. EMMA achieves promising performance on ADE 1s, 3s, and 5s. EMMA is slightly worse than Wayformer at ADE 8s, largely because the camera-only EMMA has higher depth error for longer distance than Wayformer, which uses all sensors including LiDAR, Radar, and Cameras. Note that the Wayformer baseline is based on an enhanced internal version after the publication of  (Nayakanti et al., 2023).
@@
3.1.1 Driving on the Waymo Open Motion Dataset (WOMD)
WOMD comprises 103k real-world urban and suburban driving scenarios, each lasting 20 seconds. These scenarios are further segmented into 1.1M examples, each representing a 9-second window: 1 second is used as input context, and the remaining 8 seconds serve as the prediction target. The dataset includes detailed map features such as traffic signal states and lane characteristics, along with agent states such as position, velocity, acceleration, and bounding boxes.
WOMD is typically used for the agent motion prediction task, i.e., given the agent tracks for the past 1 second on a corresponding map, predict the positions of up to 8 agents for 8 seconds into the future. We adopt the exact same setting but only train and evaluate the future motion prediction of the ego vehicle, i.e. the autonomous driving vehicle, as the planning benchmark. In addition, we leverage the ‘high-level intent command’ field, used for evaluating motion prediction, as the planning router instruction. We adapt some of the best motion prediction models, MotionLM (Seff et al., 2023) and enhanced Wayformer (Nayakanti et al., 2023), as the baselines of this planning benchmark.
During inference, sampling a final trajectory from multiple candidate trajectories plays a critical role in the final performance. Both MotionLM and Wayformer generate 192 candidate trajectories, which are subsequently aggregated into 6 clusters using k-means clustering, resulting in 6 representative trajectories. The final trajectory is selected from these 6 representative trajectories based on their associated probabilities. For fairness, we also sample multiple trajectories using a Top-$K$ decoding strategy, up to $K=24$. We then compute the pairwise L2 distance between all trajectories and select the one with the lowest average L2 distance as the final predicted trajectory, which can be viewed as the “median” trajectory among all the predictions. In the ADE calculation, we treat this median trajectory with the highest probability. All WOMD eval results are based on the validation set.
As shown in Table 1, our model achieved similar performance as the MotionLM baseline when we train only on the WOMD dataset, with Gemini pre-trained weights. When pre-trained with our internal dataset (denoted as EMMA+), our model surpasses both MotionLM and Wayformer up to ADE 5s in the future. EMMA is only slightly worse than Wayformer at ADE 8s, where camera-induced depth error becomes a prominent error resource. We note the differences in inputs between MotionLM and EMMA: MotionLM takes inputs of agent location history, agent interactions, the road graph, and traffic light states. These agent boxes are produced by specialized off-board perception models that look at both past and future observations and are trained with a large amount of carefully curated human labels, the road graph is manually generated using full run segments, and all inputs heavily use LiDAR data with superior depth estimation. In stark contrast, EMMA only takes camera images and ego vehicle history as input, without the need of any labels or additional models (besides leveraging the Gemini pre-trained weights). Notably, for this work, we added internally available image inputs, which are not currently in the released WOMD.
We further investigate the impact of the number of sampled trajectories on ADE, as illustrated in Figure 3. The results highlight that sampling from multiple trajectories, rather than a single one, leads to a notable improvement in ADE. However, the advantage of increasing the number of candidate trajectories diminishes beyond a certain point.
@@ 3.1.2 Driving on the nuScenes Dataset
The nuScenes dataset (Caesar et al., 2020) offers a comprehensive autonomous vehicle sensor suite for evaluation. It consists of 1,000 scenes, each spanning 20 seconds, and includes information from 6 cameras, 5 radars, and 1 LiDAR that collectively provide 360-degree coverage in the field of view. The dataset is fully annotated with 3D bounding boxes across 23 classes and 8 attributes. In our experiments, we follow the standard protocol for planning evaluation: predict the next 3 seconds of future driving actions based on 2 seconds of historical data. We measure the planning quality with L2 errors at 1-, 2- and 3-second time horizons, aligning with established baseline methods.
As shown in Table 2, our self-supervised EMMA achieves state-of-the-art results in planning on nuScenes, outperforming all previous supervised (with intermediate perception labels and/or human labels) and self-supervised (no extra labels) methods. Under the same self-supervised setup, EMMA outperforms previous BEV-Planner (Li et al., 2024) by 17.1% in average L2 metric; even compared to DriveVLM-Dual (Tian et al., 2024) that heavily uses intermediate perception and human labels, our self-supervised EMMA still improves the average L2 metric by 6.4%.
@@
Table 2: End-to-end motion planning experiments on nuScenes (Caesar et al., 2020). EMMA (random init) denotes models are randomly initialized; EMMA denotes models are initialized from Gemini; EMMA+ denotes models are pre-trained on our internal data. EMMA achieves state-of-the-art performance on the nuScenes planning benchmark, outperforming the supervised (with perception and/or human labels) prior art by 6.4% and self-supervised (no extra labels) prior art by 17.1%. Ego-MLP results are taken from a reproduced version in BEV-Planner.
@@ 3.2 End-to-End Motion Planning with Chain-of-Thought Reasoning on Internal Dataset
In this section, we present our studies of end-to-end planning with chain-of-thought on our internal dataset. This dataset contains millions of scenarios, orders of magnitude larger than any publicly available autonomous driving dataset. The model takes in 2 seconds of history to predict the driving actions for 5 seconds into the future.
Table 3 presents the results of our experiments on chain-of-thought reasoning applied to end-to-end planning. By adopting the chain-of-thought formulation introduced in Equation 3, we achieve a notable 6.7% improvement over the standard end-to-end planning approach detailed in Equation 2. We also conduct an ablation study to analyze the contributions of different rationale components. Our findings reveal that both driving meta-decision and critical object identification significantly enhance performance, contributing improvements of 3.0% and 1.5%, respectively. When these components are combined, the gains are even more substantial. Conversely, while scene description has a neutral impact on driving performance, it enhances the model’s explainability. These results demonstrate that chain-of-thought reasoning can meaningfully improve driving performance, particularly when its components are carefully selected and integrated.
@@
Table 3: Ablation study on chain-of-thought reasoning components. It improves end-to-end planning quality by up to 6.7% by combining all elements. In particular, driving meta-decision and critical objects contribute the improvements of 3.0% and 1.5%, respectively. The details of each component is described in Section 2.2.
@@
We also perform a series of data scaling experiments for end-to-end planning, the results of which are illustrated in Figure 4. As we train the model on a larger training set, we observe lower eval perplexities before overfitting. Our findings indicate that the driving quality of EMMA has not yet plateaued, even with the current large-scale dataset.
@@
Figure 4: EMMA data scaling experiments on our internal dataset. Each curve represents the eval perplexity for end-to-end motion planning as training proceeds with more steps. The x-axis is training compute, measured by floating-point operations (FLOPs) in log scale. The same EMMA  model is trained on four sizes of datasets that are sampled with different percentages from 3% to 100% (denoted by different colors). In general, EMMA tends to achieve better quality until overfitting when given more training compute, but it also overfits quickly on smaller datasets. We observe the driving quality has not saturated when using the full large-scale dataset.
3.3 Perception Tasks
We summarize our studies on three main perception tasks: 3D object detection, road graph estimation, and scene understanding.
3.3.1 3D Object Detection
@@ 3.3 Perception Tasks
We validate our 3D object detection performance on the 3D camera-primary detection benchmark from the Waymo Open Dataset (Sun et al., 2020) using the d Longitudinal Error Tolerant (LET) matching (Hung et al., 2024). We evaluate two versions: EMMA and EMMA+, similar to earlier sections, where EMMA+ is pre-trained on the 3D detection task using our internal dataset. The quantitative results are reported on the official test set and summarized in Figure 5.
@@
Figure 5: Camera-primary 3D object detection experiments on WOD (Sun et al., 2020) using the standard LET matching (Hung et al., 2024). EMMA+ denotes model is pre-trained on our internal data. EMMA+ achieves competitive performance on the detection benchmark in both precision/recall and F1-score metrics. Compared to state-of-the-art methods, it achieves 16.3% relative improvements in vehicle precision at the same recall or 5.5% recalls improvement at the same precision.
@@
Our findings show that after pre-training, EMMA+ achieves competitive performance on the benchmark. Since our model output set of detected boxes without individual confidence scores, we directly compare the precision/recall instead of LET-3D-AP, which is calculated based on the precision/recall curve. We also compare the commonly used F1-score, where EMMA’s F1-score is computed using the single precision/recall and other models’ F1-scores are calculated by picking the maximal F1-score on the curve (often called F1-max). Figure 5 shows the performance comparison. In generally, EMMA+ demonstrates substantial improvements over state-of-the-art methods such as BEVFormer (Li et al., 2022b), achieving a 16.3% relative increase in vehicle precision at the same recall, and a 5.5% recall improvement at the same precision. EMMA+ also achieve better F1-score than prior arts. Performance on the pedestrian class is also comparable to that of MV-FCOS3D++ (Wang et al., 2021). Additionally, we provide a performance breakdown across different ranges, highlighting that our model performs especially well in the near range. Our results underscore that with sufficient data and a large enough model, a multimodal approach can surpass specialized expert models in 3D object detection quality.
@@ 3.3.2 Road Graph Estimation
Polyline representation: dynamic sampling is better than fixed sampling. A simple polyline representation is to sample a fixed number of sparse control points per lane, e.g. two end points plus a fixed number of intermediate points to capture curvature. However, we find a better approach is to dynamically adjust the number of points per polyline according to the curvature and length of the lane. By keeping a consistent waypoint density rather than a consistent number of waypoints, we achieve a representation that more accurately captures the lane structure intricacies, yielding around a 40% to 90% difference in the metrics as shown in Figure 6.
@@ 3.4 Generalist
We explore the development of a generalist model by co-training on multiple tasks and analyzing their synergies, as summarized in Table 4. For this study, we focus on three core tasks: end-to-end planning, 3D object detection, and road graph estimation.
Co-training on all three tasks yields significant improvements, with the generalist model outperforming the single-task models by up to 5.5%. Notably, when co-training two tasks, certain combinations lead to greater gains than others. For instance, detection performance improves most when co-trained with driving, and road graph estimation similarly benefits most when paired with driving. This suggests that the driving task plays a particularly prominent and influential role, serving as a key contributor to overall performance improvements.
We attribute these results to the complementary nature of the tasks. For example, road graph estimation becomes easier when the model can accurately identify the locations of vehicles. Similarly, driving quality is closely tied to understanding agent interactions, a skill enhanced by 3D object detection. These findings suggest that pursuing a generalist model is a promising direction for future research, with the potential for deeper insights into task synergies and performance optimization.
@@
e2e planning 3D detection road graph estimation e2e planning detection road graph
@@
Table 4: Generalist co-training experiments. ($\pm$*± ∗) indicates standard deviation. By co-training on multiple tasks, EMMA gains a broader understanding of driving scenes, enabling it to handle various tasks at inference time, while enhancing individual task performance. Notably, certain task pairings yield greater benefits than others, suggesting these tasks are complementary. Co-training all three tasks together yields the best quality.
@@ 3.5 Visualizations
We present 12 diverse visual examples in Fig. 8, 9 and 10, each selected to highlight the generalizability of EMMA model across a range of scenarios. In all the scenarios, we display the model’s predictions (from left to right): end-to-end motion planning, 3D object detection, and road graph estimation.
@@
(a) A garbage bag appears on the freeway, so our predicted trajectory suggests to nudge slightly to the
right to avoid it.
@@
(b) A ladder appears on the freeway, and our predicted trajectory suggests to switch to the left lane to
bypass it appropriately.
@@
(c) we encounter a small squirrel on the road and our predicted trajectory instinctively slows down to
avoid the animal. Note EMMA wasn’t explicitly trained to detect squirrels.
@@
(d) A white dog appears in our lane, and our model predicts to slow down and yield. Our model also
accurately detects surrounding vehicles, including those in adjacent lanes and the parking lot.
@@
(e) As a construction zone blocks the left lanes, our predicted trajectory suggests passing through on
the right, while the road graph estimation correctly identifies the blocked area.
@@
(f) Our lane is blocked by construction cones, so our predicted trajectory suggests to move into the left
lane, even though it’s in the opposite direction. EMMA captured the blockage and performed a safe detour.
@@
(g) A traffic controller signals to proceed through the intersection, and our predicted trajectory aligns
with the instruction.
@@
(h) Our predicted trajectory suggests to stop as we approach an intersection with a yellow light,
demonstrating cautious and safe behavior.
@@
(i) While crossing an intersection, our predicted trajectory nudges slightly to the left due to nearby
cars and a bicyclist partially occupying our lane.
@@
(j) Our model predicts a driving trajectory to patiently wait at a red light (left). The model also
accurately predicts surrounding 3D objects (middle) and road graph with lane centers (right).
@@
(k) A fleet of fast-moving motorcyclists passes by. The predicted trajectory suggests pausing to allow
them to pass safely. Notably, motorcyclists are accurately identified by our model (middle).
@@
(l) A motorbike is moving on a narrow lane at night, and yields to the right. Our predicted trajectory
adjusts, guiding us to pass safely by nudging slightly to the left.

### v3

@@ 3.2 End-to-End Motion Planning on Public Datasets
We conduct the end-to-end planner trajectory generation experiments on two public datasets, WOMD (Chen et al., 2024a) and the nuScenes dataset (Caesar et al., 2020). EMMA is trained with the simplest end-to-end planner trajectory generation formulation as in Equation 2, unless specified otherwise. That is, given camera images, ego vehicle history, and driving intent, the model is asked to predict the future ego waypoints for a certain time horizon.
3.2.1 Driving on the Waymo Open Motion Dataset (WOMD)
Method L2 (m) 1s L2 (m) 3s L2 (m) 5s
MotionLM (Seff et al., 2023) 0.045 0.266 0.696
Wayformer (Nayakanti et al., 2023) 0.046 0.252 0.628
EMMA (based on PaLI) 0.034 0.274 0.797
EMMA+ (based on PaLI) 0.031 0.239 0.680
EMMA 0.032 0.248 0.681
EMMA (w/ CoT) 0.030 0.241 0.664
EMMA+ 0.030 0.225 0.610
EMMA+ (w/ CoT) 0.027 0.203 0.543
Table 2: End-to-end motion planning experiments on an internal planning benchmark. CoT denotes equipping with chain-of-thought reasoning (Eq. 3). EMMA+ achieves the best quality across different prediction time horizons. EMMA and EMMA+ denotes using PaLI-X (Chen et al., 2024d) as our base model, while the default EMMA and EMMA+ use Gemini as the base model. Enhanced, reproduced baselines.
For fair comparisons, we align our settings with WOMD. Additionally, we reproduce and adapt internally enhanced versions of the state-of-the-art motion prediction models, MotionLM (Seff et al., 2023) and Wayformer (Nayakanti et al., 2023), serving as our planner baselines. These baseline models are augmented with high-level command intent as an additional input.
As shown in Table 2, our model outperforms the MotionLM baseline when we train on the same dataset, with Gemini pre-trained weights. When pre-trained with our mega-scale internal dataset (denoted as EMMA+), our model outperforms both MotionLM and Wayformer. The full EFM+ (w/ CoT) surpasses the previous state-of-the-art models significantly by 13.5% at the 5s prediction horizon. We also apply the EMMA method to an open-sourced MLLM, PaLI-X (Chen et al., 2024d), denoted as EMMA (PaLI). We show that EMMA can generalize well across different MLLMs with a large amount of training data, yielding better quality (at 1s and 3s) than previous state-of-the-art baselines.
We note the differences in inputs between MotionLM and EMMA: MotionLM takes inputs of agent location history, agent interactions, the road graph, and traffic light states. These agent boxes are produced by specialized off-board perception models that look at both past and future observations and are trained with a large amount of carefully curated human labels, the road graph is manually generated using full run segments, and all inputs heavily use LiDAR data with superior depth estimation. In stark contrast, EMMA only takes camera images and ego vehicle history as input, without the need of any labels or additional models (besides leveraging the Gemini pre-trained weights).
During inference, sampling a final trajectory from multiple candidates plays a critical role in the final performance. Both MotionLM and Wayformer generate 192 candidate trajectories, which are subsequently aggregated into 6 clusters using k-means clustering, resulting in 6 representative trajectories to be selected as the final output according to their probabilities. For fairness, we also sample multiple trajectories using a Top-$K$ decoding strategy, up to $K=24$. We then compute the pairwise L2 distance between all trajectories and select the one with the lowest average L2 distance as the final predicted trajectory, which can be viewed as the “median” trajectory among all the predictions. We investigate the impact of the number of sampled trajectories on ADE, as illustrated in Figure 3. The results highlight that sampling from multiple trajectories leads to a notable improvement in ADE, however, with diminishing return.
@@ 3.2.2 Driving on the nuScenes Dataset
In our experiments, we follow the standard protocol of nuScenes for planning evaluation: predict the next 3 seconds of future driving actions based on 2 seconds of historical data. We measure the planning quality with L2 errors at 1-, 2- and 3-second time horizons, aligning with established baseline methods, in particular BEV-Planner (Li et al., 2024).
We train and evaluate EMMA with the simplest end-to-end planner trajectory generation formulation as in Equation 2 (self-supervised, without chain-of-thought reasoning nor generalist training). As shown in Table 3, our self-supervised EMMA achieves state-of-the-art results in planning on nuScenes, outperforming all previous supervised (with intermediate perception labels and/or human labels) and self-supervised (no extra labels) methods. Under the same self-supervised setup, EMMA outperforms BEV-Planner (Li et al., 2024) by 17.1% in average L2 metric; even compared to OmniDrive (Wang et al., 2024a) that heavily uses intermediate perception human labels, our self-supervised EMMA improves the average L2 metric by 12.1%.
@@
Table 3: End-to-end motion planning experiments on nuScenes (Caesar et al., 2020). EMMA (random init) denotes models are randomly initialized; EMMA denotes models are initialized from Gemini; EMMA+ denotes models that are pre-trained on our mega-scale internal data. EMMA achieves state-of-the-art performance on the nuScenes planning benchmark, outperforming the supervised (with perception and/or human labels) prior art by 6.4% and self-supervised (no extra labels) prior art by 17.1%. Ego-MLP results are taken from a reproduced version in BEV-Planner.
@@ 3.3 End-to-End Motion Planning with Chain-of-Thought Reasoning on Internal Dataset
In this section, we present our studies of end-to-end planning with chain-of-thought on our internal dataset. This dataset contains 24 millions of scenarios, orders of magnitude larger than any publicly available autonomous driving dataset. The model takes in 2 seconds of history to predict the driving actions for 5 seconds into the future.
Table 4 presents the results of our experiments on chain-of-thought reasoning applied to end-to-end planning. By adopting the chain-of-thought formulation (Equation 3), we achieve a notable 6.7% improvement over the standard end-to-end planning approach detailed in Equation 2. We also conduct an ablation study to analyze the contributions of different rationale components. Our findings reveal that both driving meta-decision and critical object identification significantly enhance performance, contributing improvements of 3.0% and 1.5%, respectively. When these components are combined, the gains are even more substantial. Conversely, while scene description has a neutral impact on driving performance, it enhances the model’s explainability. These results demonstrate that chain-of-thought reasoning can meaningfully improve driving performance, particularly when its components are carefully selected and integrated.
@@
Table 4: Ablation study on chain-of-thought reasoning components. It improves end-to-end planning quality by up to 6.7% by combining all elements. In particular, driving meta-decision and critical objects contribute the improvements of 3.0% and 1.5%, respectively. The details of each component is described in Section $2.$
@@
We also perform a series of data scaling experiments for end-to-end planning, the results of which are illustrated in Figure 4. As we train the model on a larger training set, we observe lower eval perplexities before overfitting. Our findings indicate that the driving quality of EMMA has not yet plateaued, even with the current mega-scale dataset.
@@
Figure 4: EMMA data scaling experiments on our mega-scale internal dataset. Each curve represents the eval perplexity for end-to-end motion planning as training proceeds with more steps. The x-axis is training compute, measured by floating-point operations (FLOPs) in log scale. The same EMMA model is trained on four sizes of datasets that are sampled with different percentages from 3% to 100% (denoted by different colors). In general, EMMA tends to achieve better quality until overfitting when given more training compute, but it overfits quickly on smaller datasets. We observe the driving quality has not saturated when using the full large-scale dataset.
3.4 3D Object Detection
@@
We validate our 3D object detection performance on the 3D camera-primary detection benchmark from the Waymo Open Dataset (Sun et al., 2020) using the Longitudinal Error Tolerant (LET) matching (Hung et al., 2024). We evaluate two versions: EMMA and EMMA+, similar to earlier sections, where EMMA+ is pre-trained on the 3D detection task using our internal dataset. The quantitative results are reported on the official test set and summarized in Figure 5.
@@
Figure 5: Camera-primary 3D object detection experiments on WOD (Sun et al., 2020) using the standard LET matching (Hung et al., 2024). EMMA+ achieves competitive performance on the detection benchmark in both precision/recall and F1-score metrics. Compared to state-of-the-art methods, it achieves 16.3% relative improvements in vehicle precision at the same recall or 5.5% recall improvement at the same precision.
@@
Our findings show that after pre-training, EMMA+ achieves competitive performance on the benchmark. Since our model produces a set of detected boxes without individual confidence scores, we compare the precision/recall instead of LET-3D-AP, which is calculated based on the precision/recall curve. We also compare the commonly used F1-score, where EMMA’s F1-score is computed using the single precision/recall and other models’ F1-scores are calculated by picking the maximal F1-score on the curve (often called F1-max).
Figure 5 shows the performance comparison. In generally, EMMA+ demonstrates substantial improvements over state-of-the-art methods such as BEVFormer (Li et al., 2022b), achieving a 16.3% relative increase in vehicle precision at the same recall, and a 5.5% recall improvement at the same precision. EMMA+ also achieve better F1-score than prior arts. Performance on the pedestrian class is also comparable to that of MV-FCOS3D++ (Wang et al., 2021). Additionally, we provide a performance breakdown across different ranges, highlighting that our model performs especially well in the near range. Our results underscore that with sufficient data and a large enough model, a multimodal approach can surpass specialized expert models in 3D detection quality.
@@ 3.4.1 Road Graph Estimation
Polyline representation: dynamic sampling is better than fixed sampling. A simple polyline representation is to sample a fixed number of sparse control points per lane, e.g. two end points plus a fixed number of intermediate points to capture curvature. However, we find a better approach is to dynamically adjust the number of points per polyline according to the curvature and length of the lane. By keeping a consistent waypoint density rather than a consistent number of waypoints, we achieve a representation that more accurately captures the lane structure intricacies, yielding around a 40% to 90% difference in the metrics as shown in Figure 6. By adapting the waypoint density to the road geometry, particularly in areas with sharper curves or varying lane lengths, we achieve a flexible representation that more accurately captures the lane structure intricacies
@@ 3.5 Generalist
We explore the development of the EMMA Generalist by co-training on multiple tasks and analyzing their synergies, as summarized in Table 5. For this study, we focus on three core tasks: end-to-end planning, 3D object detection, and road graph estimation.
Co-training on all three tasks yields significant improvements, with the generalist model outperforming the single-task models by up to 5.5%. We attribute these results to the complementary nature of the tasks. For example, road graph estimation becomes easier when the model can accurately identify the locations of vehicles. Similarly, driving quality is closely tied to understanding agent interactions, a skill enhanced by 3D object detection. Paring the mixture down to only two tasks still yields improvements, with certain combinations leading to greater gains than others. For instance, detection performance improves most when co-trained with driving, and road graph estimation similarly benefits most when paired with driving. This suggests the driving task plays a prominent and influential role, serving as a key contributor to overall performance improvements.
These findings suggest that pursuing a generalist model is a promising direction for future research, with the potential for deeper insights into task synergies and performance optimization.
@@
e2e planning 3D detection road graph e2e planning detection road graph
@@
Table 5: Generalist co-training experiments. ($\pm*$) indicates standard deviation. By co-training on multiple tasks, EMMA gains a broader understanding of driving scenes, enabling it to handle various tasks at inference time, while enhancing individual task performance. Notably, certain task pairings yield greater benefits than others, suggesting these tasks are complementary. Co-training all three tasks together yields the best quality.
@@ 3.6 Visualizations
(a) A garbage bag appears on the freeway, so our predicted trajectory suggests to nudge slightly to the right to avoid it.
@@
(b) A ladder appears on the freeway, and our predicted trajectory suggests to switch to the left lane to bypass it appropriately.
@@
(c) we encounter a small squirrel on the road and our predicted trajectory instinctively slows down to avoid the animal. Note EMMA wasn’t explicitly trained to detect squirrels.
@@
(d) A white dog appears in our lane, and our model predicts to slow down and yield. Our model also accurately detects surrounding vehicles, including those in adjacent lanes and the parking lot.
@@
(e) As a construction zone blocks the left lanes, our predicted trajectory suggests passing through on the right, while the road graph estimation correctly identifies the blocked area.
@@
(f) Our lane is blocked by construction cones, so our predicted trajectory suggests to move into the left lane, even though it’s in the opposite direction. EMMA captured the blockage and performed a detour.
@@
(g) A traffic controller signals to proceed through the intersection, and our predicted trajectory aligns with the instruction.
@@
(h) Our predicted trajectory suggests to stop as we approach an intersection with a yellow light, demonstrating cautious and safe behavior.
@@
(i) While crossing an intersection, our predicted trajectory nudges slightly to the left due to nearby cars and a bicyclist partially occupying our lane.
@@
(j) Our model predicts a driving trajectory to patiently wait at a red light (left). The model also accurately predicts surrounding 3D objects (middle) and road graph with lane centers (right).
@@
(k) A fleet of fast-moving motorcyclists pass by. The predicted trajectory suggests pausing to allow them to pass safely. Notably, motorcyclists are accurately identified by our model (middle).
@@
(l) A motorbike is moving on a narrow lane at night, and yields to the right. Our predicted trajectory adjusts, guiding us to pass safely by nudging slightly to the left.